# Learning Generalizable Tool-use Skills through Trajectory Generation

### v4

@@ I INTRODUCTION
Building autonomous systems that leverage tools can greatly enhance efficiency and assist humans in completing many common tasks in everyday life [1, 2, 3, 4, 5, 6, 7]. As humans, we possess an innate ability to adapt quickly to use novel tools. However, replicating such adaptability in autonomous systems presents a significant challenge. To solve this task of novel tool manipulation, prior work has explored different representations for tools. A good tool representation should contain a rich visual understanding of the object and be useful for downstream physical interactions. Prior work [2],[8] use data-driven approaches to learn the latent representations for tools but such representation cannot generalize because of the lack of compositionality and interpretability. Another line of the work studies keypoints as a representation for tool which works only for rigid object manipulation including hammering, pushing and reaching.
@@
Our goal is to train a policy to solve various manipulation tasks with multiple tools, including tools that were not seen during training. We propose a novel approach, ToolGen, which learns tool-use skills via trajectory generation and sequential pose optimization. Given the scene, the goal, and a tool, ToolGen first generates a point cloud of a tool in the desired initial pose, and it subsequently predicts how this generated tool would move to perform the task. Finally, we sequentially align the actual tool to the generated tool to extract the actions for the agent to execute. We evaluate ToolGen against several baselines in deformable object manipulation with diverse tasks, goals, and tools. Impressively, with just a single model trained across all tasks and tools, ToolGen significantly outperforms the baselines and generalizes to many novel tools. Further, ToolGen achieves this despite being trained on demonstrations from just one tool for each task.
To summarize our contribution, we propose ToolGen, which represents tool use via trajectory generation. We have shown that generating a point cloud trajectory of the tool can effectively capture the essence of tool use, i.e. how the tool should be placed in relation to the dough and how it should move over time, which allows us to generalize to a variety of unseen tools and goals. Furthermore, we use sim-to-real transfer to demonstrate our method’s effectiveness on three real world manipulation tasks with unseen novel tools and different goals.
@@ II Related Work
Learning Generalizable Tool-use Skills: Prior work has explored training robots to perform manipulation tasks with tools. To enable generalization, some previous approaches predict intermediate “affordances” and then generate actions based on these affordances [2, 9]. For example, affordances like grasping points or functional points and be represented as key points [2, 3, 4]. Similarly, concepts like contacts and forces [10, 11] can also be used. However, obtaining labels for these affordances can be difficult, and such affordance labels do not easily extend to more complex tasks such as deformable object manipulation as we explore in this work. Another approach is to discover affordance regions in a self-supervised way by running parameterized motion primitives [2] or affordance-conditioned policies [3, 4] in simulation. In the image space, prior works have explored training an action-conditioned video prediction model [1] for planning actions for different tools. However, the video prediction model lacks 3D structure and has difficulty representing fine-grained action trajectories. Another research direction for generalizable tool use is to utilize the pretrained Large Language Models (LLMs) for long horizon reasoning. Prior work[12] designed four consecutive modules to prompt LLMs to directly generate code for robotic tool use. However, they make an assumption that we have state information of the tools and objects and directly include them into the prompt for LLMs. For deformable objects, state estimation is very challenging so this approach doesn’t generalize to deformable object manipulation with tool use.
@@ III Problem statement and assumptions
Consider a set of point clouds $(P^{o},P^{g},P^{tool})$, where $P^{o}$ represents the initial observation of the scene, $P^{g}$ stands for the goal, and $P^{tool}$ for a tool to use for execution. Our task is to predict an actions sequence of horizon $H$, where the tool transforms the initial pose into a predicted target pose. The actions is represented by a transformations sequence $T_{0:H}$.
@@ IV-A Representing tool-use through point cloud trajectory generation
$G_{reset}$ is a PointFlow-based [19] encode-decoder generation model. It conditions on the point cloud of the existing tool $P^{tool}$, the initial scene observation $P^{o}$, and the goal $P^{g}$, to reconstruct the tool at “reset pose” $P^{gen}_{0}$. The architecture of our PointFlow-based [19] generator $G_{reset}$ is shown in Fig. 2 (b) (top). It encodes the tool points and the concatenation initial and target states to two sets of latent features with separate PointNet++ [20] encoders. These latent features are concatenated and inputted through an MLP to produce an estimation of Gaussian distribution. We then take a sample from this estimated distribution as the input of a PointFlow [19] decoder, which outputs the reconstruction point cloud of the given tool at the reset pose $P^{gen}_{0}$.
The second part $G_{path}$ works on predicting a sequence of transformations of how this generated tool would move to achieve the task. The architecture of the path generator is shown in Fig. 2 (a) (top right). We follow the design in ToolFlowNet [21] to train a policy model through Behavior Cloning, which optimizes a combined loss of point content loss and consistency loss. The $P^{gen}_{0}$ from $G_{reset}$ is concatenated together with the initial scene observation $P^{o}$, and the goal state $P^{g}$ and fed into the model. Transformations of $H-1$ time-steps, $T^{gen}_{1:H}$, are generated for the tool. Details for $G_{path}$ can be found in Appendix A-B.
@@ IV-B Execution via sequential pose optimization
1:Input: The current observation of the dough $P^{o}$, the existing tool $P_{tool}$, and the point cloud trajectory for the generated tool $P_{gen}^{o:H}$
@@ IV-C Implementation details
We train a single set of modules ($G_{reset},G_{path}$) across a compact demonstration dataset comprised of multiple tasks rather than training separate networks for each task. To achieve this, we introduces a scoring module $D_{score}$ to evaluate and select tool for each task. Details of $D_{score}$ module is shown in Appendix A-A. In training data, we collect 200 demonstration trajectories for each task performed with just one training tool. Despite the limited training data, our model is demonstrated to be capable of generalizing to various unseen tools in both simulation and real world. See Appendix B-A for more information on our demonstration dataset.
In trajectory optimization, we use the quaternion representation for the orientation of the transformation, and project the values onto a unit ball after each gradient update. Here, we use a step size of $10^{-2}$, and $\lambda_{c}=0.1$. For optimizing the delta poses, we use the 3-DoF Euler angles representation with a step size of $10^{-3}$, a regularization factor of $\lambda_{r}=0.1$, and we use the euclidean norm to regularize the translation as well as the rotation.
@@ V-A Tasks and baselines
Baselines: We evaluate the following baselines with different action representations. All of the baselines regress to reset transformations and delta poses, except for BC-E2E which predicts delta poses directly from the initial configuration without a reset transformation. Details on the architectures of the baselines are described in Appendix B-B.
@@
  • TFN-Traj. Behavioral cloning that regresses to the reset transformation, moves the tool to the predicted reset pose, and then uses the updated scene to predict subsequent delta poses with the ToolFlowNet-based [21] trajectory model described in Appendix A-B.
We examine three settings, each presenting a greater level of difficulty, detailed in Sec. V-B, Sec. V-C, and Sec. V-D, respectively. We demonstrate that ToolGen is robust to these generalization challenges and maintains superior performance over the baselines. We additionally conduct ablation studies by removing the path generator of ToolGen, detailed in Appendix C-A.
@@ V-B Leveraging training tools at test time
Figure 6: Results of ToolGen on real world observations for Cut (top) and Rol (middle) and Scoop(bottom). For each task, we visualize initial dough observation, the goal, the real world tool, the policy rollout of our trained policy and the final state of the policy. As a result, ToolGen can effectively generate manipulation trajectories from real world observations even though the model is trained entirely in simulation.
@@ V-D Generalization to unseen tools
Finally, we evaluate the methods on a far more challenging scenario, in which our agents are given unseen tools. We evaluate each novel tool on 10 held out goals for each task and average their performances. See Fig. 3 for a visualization of the novel tools we consider. Since the novel tools are also in arbitrary initial poses, this scenario requires the method to be robust to tool shapes as well as initial poses of the tool. Fig. 4(a) and Table I shows the quantitative results of all the methods, and Fig. 5 show examples of rollouts by ToolGen (ours) and the baseline TFN-Traj. All of the baselines fail to obtain a high performance, especially in the more challenging task of scooping (see Table I). In contrast, ToolGen can leverage completely unseen tools in meaningful ways. This is because ToolGen leverages trajectory generation to alleviate the issues of distribution shift. It further uses a non-learned optimization procedure (gradient descent with multiple random initializations), which also does not suffer from a distribution shift. For more analysis, please see our Appendix C-A.
@@ V-E Inference on real world observations
For our real world experiments, we select three representative tasks, Cut, Roll and Scoop(Large) to test our trained policy. In each task, we select two real world tools and attach each tool to a mount so that it can stay on the tool hanger for the robot to pick up. Details of our environment setup can be seen in Figure 7 and Appendix C-D. Our ToolGen model is trained entirely with simulation data. To demonstrate the robustness of ToolGen, we record point clouds of tool and dough from the real world and use ToolGen to predict the trajectory of the real world tool. To obtain the point clouds from the real world, we use three Azure Kinect cameras to record the initial dough and the tool point clouds. We then manually manipulate the dough to a desired shape and record the final point cloud as the goal point cloud. We record the point cloud of the real dough at its initial and goal states and concatenate them with the tool point cloud. The initial pose of the tool is entailed in the point cloud input. The model then output a trajectory of horizon $H=50$. In the execution stage, we use a mold to restore the dough to the recorded initial state. The robot picks up the tool and move to the recorded initial pose and executes the trajectory, where the tool first transforms from the initial pose to reset pose, and then moves through to execute the produced trajectory on the real dough. Fig. 6 includes qualitative results of the robot executing the policy’s rollouts. For each of the three different tasks, we show an example of the robot using one of the test tools to reshape the dough. Quantitatively, despite the gap in point cloud observations between sim and real, our method can effectively generalize to unseen tool in the real world with an average normalized decrease of Chamfer Distance of 0.77 as shown in Table II. This metric shows how close the final state gets to the goal compared to the initial state of the dough. The larger normalized decrease of Chamfer Distance indicates better performance. Compared to the baseline method BC-Latent, our method ToolGen outperforms it on all three tasks by a large margin. BC-Latent fails to generalize to some unseen real world tools and generates transformations that oscillate without further movements for those tools. In contrast, our method successfully generalizes to all tasks with different goals and tools. Hence, our method demonstrates smaller performance variance. To prove that our policy is comparable to human performance, two volunteers are asked to perform the same manipulation tasks as that for robots. From the table we can see that our policy’s task performance is very close to humans with the largest difference of 0.08 in normalized decrease of Chamfer Distance. We also notice that Scoop generally has worst performance compared to other two tasks because the real world dough we are using is so sticky that both the human and our trained policy struggle with detaching the scooped piece from the whole dough.
@@ VI Conclusion and limitations
Limitations: Our method has several limitations: First, our method’s execution time is considerably longer compared to that of a trained policy, due to the time needed for generating point clouds and optimizing the current tool’s poses. We anticipate that the use of faster techniques for sequential pose optimization, such as second-order methods, could speed up our method. Secondly, as our point cloud generator is trained on very limited tools, it is sometimes unable to generate accurate point clouds for novel tools and thus the alignment process could fail. A promising direction is to train on more variations of the tool to improve the generation process and make alignment easier. Further details on these failure cases can be found in Appendix C-B.
@@ B-B Details on baselines
  • TFN-Traj. Behavioral cloning that regresses to the reset transformation, moves the tool to the predicted reset pose, and then uses the updated scene to predict subsequent delta poses for the tool with the ToolFlowNet-based [21] trajectory model described in Appendix A-B.
@@ C-A Ablation studies
C-B Failure cases
@@ C-C Effects of hyperparameters
TABLE V: The effects of hyperparameters in sequential pose estimation. Each entry shows the performance cross all tasks with a particular combination of hyperparameters.
To investigate how sensitive ToolGen is to hyperparameters during sequential pose estimation, we vary $\lambda_{c}$ and $\lambda_{r}$ when optimizing for the transformations. To eliminate any stochasticity form the generation process, we only generate the trajectories once and use the same set of generated trajectories for optimization. Table V shows the performances of executing the trajectories that are optimized with different hyperparameters. In summary, $\lambda_{c}$, which penalizes collisions between the tool in reset pose and the environment, seems to have a greater effect than $\lambda_{r}$. This is because the alignment at the rest pose is being used to optimize the subsequent poses, and the error in optimizing the reset pose might cascade to the later process. We also observe that choosing a larger $\lambda_{r}$ will decrease the variance in performance. This is because a larger $\lambda_{r}$ will penalize large motions and encourage smaller and safer motions. It is worth noting that even with a large variation of these hyperparameters, ToolGen still almost always outperforms the baselines.
C-D Real World Experiment Details
C-D1 Environment setup

### v5

@@ I INTRODUCTION
Building autonomous systems that leverage tools can greatly enhance efficiency and assist humans in completing many common tasks in everyday life [1, 2, 3, 4, 5, 6, 7]. As humans, we possess an innate ability to adapt quickly to use novel tools. However, replicating such adaptability in autonomous systems presents a significant challenge. To solve this task of novel tool manipulation, prior work has explored different representations for tools. A good tool representation should contain a rich visual understanding of the object and be useful for downstream physical interactions. Prior work [2] uses data-driven approaches to learn the latent representations for tools but such representation cannot generalize because of the lack of compositionality and interpretability. Another line of the work studies keypoints as a representation for tool which works only for rigid object manipulation including hammering, pushing and reaching.
@@
Our goal is to train a policy to solve various manipulation tasks with multiple tools, including tools that were not seen during training. We propose a novel approach, ToolGen, which learns tool-use skills via trajectory generation and sequential pose optimization. Given the scene, the goal, and a tool, ToolGen first generates a point cloud of a tool in the desired initial pose, and it subsequently predicts how this generated tool would move to perform the task. Finally, we sequentially align the actual tool to the generated tool to extract the actions for the agent to execute. Fig 1 offers an overview of our task setting and ToolGen’s outputs. We evaluate ToolGen against several baselines in deformable object manipulation with diverse tasks, goals, and tools. Impressively, with just a single model trained across all tasks and tools, ToolGen significantly outperforms the baselines and generalizes to many novel tools. Further, ToolGen achieves this despite being trained on demonstrations from just one tool for each task.
To summarize our contribution, we propose ToolGen, which represents tool use via trajectory generation. We have shown that generating a point cloud trajectory of the tool can effectively capture the essence of tool use, i.e. how the tool should be placed in relation to the dough and how it should move over time, which allows us to generalize to a variety of unseen tools and goals. Furthermore, we transfer the policy to the real world without any finetuning to demonstrate our method’s effectiveness on three real world manipulation tasks with unseen novel tools and different goals.
@@ II Related Work
Learning Generalizable Tool-use Skills: Prior works have explored training robots to perform manipulation tasks with tools. To enable generalization, some approaches predict intermediate “affordances” and then generate actions based on these affordances [2, 8, 9, 10]. For example, affordances like grasping points or functional points and be represented as key points [2, 3, 4, 9, 10]. Similarly, concepts like contacts and forces [11, 12] can also be used. However, obtaining labels for these affordances can be difficult, and such affordance labels do not easily extend to deformable object manipulation, since the contacts with deformable objects (e.g. rolling a piece of dough) are complex and cannot be modeled by a few keypoints. Comparing to these methods, our method is capable of learning from unlabeled interaction data, as it implicitly learns affordances from the point clouds of the tool and the dough. This data-driven approach is similar to prior work [13], but we do not explicitly specify the structure of the shape embedding space, leaving more flexibility in tool shapes.
Another approach is to discover affordance regions in a self-supervised way by running parameterized motion primitives [2] or affordance-conditioned policies [3, 4] in simulation. In the image space, prior works have explored training an action-conditioned video prediction model [1] for planning actions for different tools. However, the video prediction model lacks 3D structure and has difficulty representing fine-grained action trajectories. Another research direction for generalizable tool use is to utilize the pretrained Large Language Models (LLMs) for long horizon reasoning. Prior work [14] designs four consecutive modules to prompt LLMs to directly generate code for robotic tool use. However, they make an assumption that we have state information of the tools and objects and directly include them into the prompt for LLMs. For deformable objects, state estimation is very challenging so this approach doesn’t generalize to deformable object manipulation with tool use.
@@ III Problem statement and assumptions
Consider a set of point clouds $(P^{o},P^{g},P^{tool})$, where $P^{o}$ represents the initial observation of the scene, $P^{g}$ stands for the goal, and $P^{tool}$ for a tool to use for execution. Our task is to predict an actions sequence of horizon $H$, where the tool transforms the initial pose into a predicted target pose. The actions is represented by a transformations sequence $T_{0:H}$. Here, all the point cloud positions as well as the objects’ orientations are relative to a reference frame located at the dough center. This design allows us to perform manipulation that is agnostic to the location of the dough on the table.
@@ IV-A Representing tool-use through point cloud trajectory generation
$G_{reset}$ is a PointFlow-based [21] encode-decoder generation model. It conditions on the point cloud of the existing tool $P^{tool}$, the initial scene observation $P^{o}$, and the goal $P^{g}$, to reconstruct the tool at “reset pose”, $P^{gen}_{0}$. The architecture of our PointFlow-based [21] generator $G_{reset}$ is shown in Fig. 2 (a) (top). It encodes the tool points and the concatenation of initial and target dough points to two sets of latent features with separate PointNet++ [22] encoders. These latent features are concatenated and inputted through an MLP to produce an estimation of Gaussian distribution. We then take a sample from this estimated distribution as the input of a PointFlow [21] decoder, which outputs the reconstruction point cloud of the given tool at the reset pose $P^{gen}_{0}$.
The second part $G_{path}$ works on predicting a sequence of transformations of how this generated tool would move to achieve the task. The architecture of the path generator is shown in Fig. 2 (a) (top right). We follow the design in ToolFlowNet [23] to train a policy model through Behavior Cloning, which optimizes a combined loss of point content loss and consistency loss. The $P^{gen}_{0}$ from $G_{reset}$ is concatenated together with the initial scene observation $P^{o}$, and the goal state $P^{g}$ and passed into the model. Transformations of $H-1$ time-steps, $T^{gen}_{1:H}$, are generated for the tool. Details for $G_{path}$ can be found in Appendix A-B on the website.
@@ IV-B Execution via sequential pose optimization
1:Input: The current observation of the dough $P^{o}$, the existing tool $P^{tool}$, and the point cloud trajectory for the generated tool $P^{gen}_{0:H}$
@@ IV-C Implementation details
We train a single set of modules ($G_{reset},G_{path}$) across a compact demonstration dataset comprised of multiple tasks rather than training separate networks for each task. To achieve this, we introduces a scoring module $D_{score}$ to evaluate and select tool for each task. Details of $D_{score}$ module is shown in Appendix A-A on the website. In training data, we collect 200 demonstration trajectories for each task performed with just one training tool. Despite the limited training data, our model is demonstrated to be capable of generalizing to various unseen tools in both simulation and real world. See Appendix B-A on the website for more information on our demonstration dataset.
In trajectory optimization, we use the quaternion representation for the orientation of the transformation, and project the values onto a unit ball after each gradient update. Here, we use a step size of $10^{-2}$, and $\lambda_{c}=0.1$. For optimizing the delta poses, we use the 3-DoF Euler angles representation with a step size of $10^{-3}$, a regularization factor of $\lambda_{r}=0.1$, and we use the euclidean norm to regularize the translation as well as the rotation. We use a greedy IK solver [24] to obtain the robot actions from the simulation and the real world, and we find it to work well in our tasks. One could also use IK as the main objective in the sequential pose optimization step to produce better poses for IK to solve.
@@ V-A Tasks and baselines
Baselines: We evaluate the following baselines with different action representations. All of the baselines regress to reset transformations and delta poses, except for BC-E2E which predicts delta poses directly from the initial configuration without a reset transformation. Details on the architectures of the baselines are described in Appendix B-B on the website.
  • TrajOpt Oracle. Differentiable trajectory optimization with ground truth dynamics from the simulator.
@@
  • TFN-Traj. Behavioral cloning that regresses to the reset transformation, moves the tool to the predicted reset pose, and then uses the updated scene to predict subsequent delta poses with the ToolFlowNet-based [23] trajectory model described in Appendix A-B on the website.
We examine three settings, each presenting a greater level of difficulty, detailed in Sec. V-B, Sec. V-C, and Sec. V-D, respectively. We demonstrate that ToolGen is robust to these generalization challenges and maintains superior performance over the baselines. We additionally conduct ablation studies by removing the path generator of ToolGen, detailed in Appendix C-A on the website.
@@ V-B Leveraging training tools at test time
Figure 6: Results of ToolGen on real world observations for Cut (top) and Roll (middle) and Scoop(bottom). For each task, we visualize initial dough observation, the goal, the real world tool, the policy rollout of our trained policy and the final state of the policy. As a result, ToolGen can effectively generate manipulation trajectories from real world observations even though the model is trained entirely in simulation.
@@ V-D Generalization to unseen tools
Finally, we evaluate the methods on a far more challenging scenario, in which our agents are given unseen tools. We evaluate each novel tool on 10 held out goals for each task and average their performances. See Fig. 3 for a visualization of the novel tools we consider. Since the novel tools are also in arbitrary initial poses, this scenario requires the method to be robust to tool shapes as well as initial poses of the tool. Fig. 4(a) and Table I shows the quantitative results of all the methods, and Fig. 5 show examples of rollouts by ToolGen (ours) and the baseline TFN-Traj. All of the baselines fail to obtain a high performance, especially in the more challenging task of scooping (see Table I). In contrast, ToolGen can leverage completely unseen tools in meaningful ways. This is because ToolGen leverages trajectory generation to alleviate the issues of distribution shift. It further uses a non-learned optimization procedure (gradient descent with multiple random initializations), which also does not suffer from a distribution shift. For more analysis, please see our Appendix C-A on the website.
@@ V-E Inference on real world observations
For our real world experiments, we select three representative tasks, Cut, Roll and Scoop(Large) to test our trained policy. In each task, we select two real world tools and attach each tool to a mount so that it can stay on the tool hanger for the robot to pick up. Details of our environment setup can be seen in Figure 7 and Appendix C-E on the website. Our ToolGen model is trained entirely with simulation data. To demonstrate the robustness of ToolGen, we record point clouds of tool and dough from the real world and use ToolGen to predict the trajectory of the real world tool. To obtain the point clouds from the real world, we use three Azure Kinect cameras to record the initial dough and the tool point clouds. We then manually manipulate the dough to a desired shape and record the final point cloud as the goal point cloud. We record the point cloud of the real dough at its initial and goal states and concatenate them with the tool point cloud. The initial pose of the tool is entailed in the point cloud input. The model then output a trajectory of horizon $H=50$. In the execution stage, we use a mold to restore the dough to the recorded initial state. The robot picks up the tool and move to the recorded initial pose and executes the trajectory, where the tool first transforms from the initial pose to reset pose, and then moves through to execute the produced trajectory on the real dough. Fig. 6 includes qualitative results of the robot executing the policy’s rollouts. For each of the three different tasks, we show an example of the robot using one of the test tools to reshape the dough. Quantitatively, despite the gap in point cloud observations between sim and real, our method can effectively generalize to unseen tool in the real world with an average normalized decrease of Chamfer Distance of 0.77 as shown in Table II. This metric shows how close the final state gets to the goal compared to the initial state of the dough. The larger normalized decrease of Chamfer Distance indicates better performance. Compared to the baseline method BC-Latent, our method ToolGen outperforms it on all three tasks by a large margin. BC-Latent fails to generalize to some unseen real world tools and generates transformations that oscillate without further movements for those tools. In contrast, our method successfully generalizes to all tasks with different goals and tools. Hence, our method demonstrates smaller performance variance. To prove that our policy is comparable to human performance, two volunteers are asked to perform the same manipulation tasks as that for robots. From the table we can see that our policy’s task performance is very close to humans with the largest difference of 0.08 in normalized decrease of Chamfer Distance. We also notice that Scoop generally has worst performance compared to other two tasks because the real world dough we are using is so sticky that both the human and our trained policy struggle with detaching the scooped piece from the whole dough.
@@ VI Conclusion and limitations
Limitations: Our method has several limitations: First, our method’s execution time is considerably longer compared to that of a trained policy, due to the time needed for generating point clouds and optimizing the current tool’s poses. Quantitative results are shown in Appendix C-B on the website. We anticipate that the use of faster techniques for sequential pose optimization, such as second-order methods, could speed up our method. Secondly, as our point cloud generator is trained on limited tools, it is sometimes unable to generate accurate point clouds for novel tools and thus the alignment process could fail. A promising direction is to train on more variations of the tool to improve the generation process and make alignment easier. Further details on these failure cases are shown in Appendix C-C on the website.
@@ B-B Details on baselines
  • TFN-Traj. Behavioral cloning that regresses to the reset transformation, moves the tool to the predicted reset pose, and then uses the updated scene to predict subsequent delta poses for the tool with the ToolFlowNet-based [23] trajectory model described in Appendix A-B on the webiste.
@@ C-A Ablation studies
Method Average Inference Time Average Execution Time
TFN-Traj $0.2$s $23.0$s
ToolGen (Ours) $22.7$s $19.1$s
TABLE V: Execution times averaged for all simulation tasks.
@@
C-B Execution time
We further compare the average inference time of ToolGen and a baseline in Table V. Due to the sequential pose optimization step, ToolGen requires significantly more time during inference compared to its baseline, TFN-Traj, which only requires a single forward pass in the networks. We leave improving the time efficiency of ToolGen’s trajectory generation for future work.
C-C Failure cases
@@ C-D Effects of hyperparameters
TABLE VI: The effects of hyperparameters in sequential pose estimation. Each entry shows the performance cross all tasks with a particular combination of hyperparameters.
To investigate how sensitive ToolGen is to hyperparameters during sequential pose estimation, we vary $\lambda_{c}$ and $\lambda_{r}$ when optimizing for the transformations. To eliminate any stochasticity form the generation process, we only generate the trajectories once and use the same set of generated trajectories for optimization. Table VI shows the performances of executing the trajectories that are optimized with different hyperparameters. In summary, $\lambda_{c}$, which penalizes collisions between the tool in reset pose and the environment, seems to have a greater effect than $\lambda_{r}$. This is because the alignment at the rest pose is being used to optimize the subsequent poses, and the error in optimizing the reset pose might cascade to the later process. We also observe that choosing a larger $\lambda_{r}$ will decrease the variance in performance. This is because a larger $\lambda_{r}$ will penalize large motions and encourage smaller and safer motions. It is worth noting that even with a large variation of these hyperparameters, ToolGen still almost always outperforms the baselines.
C-E Real World Experiment Details
C-E1 Environment setup